Circular No. 7322

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

BL LACERTAE F. Montagni, E. Massaro, and R. Nesci, Istituto Astronomico, Univer-

sitá ‘La Sapienza’, Rome; G. Tagliaferri, Brera Astronomical Observatory; and G. Tosti, Astronomical Observatory, Perugia University, report that BL Lac is again at a brightness level comparable to the maximum reached during the large 1997 outburst: Nov. 27.75 UT, V = 13.4 ± 0.03; 28.75, 13.0 ± 0.03. Further multifrequency observations are encouraged. A BeppoSAX observation of BL Lac is scheduled between Dec. 5.455 and 6.729.

SUPERNOVA 1998S IN NGC 3877 S. D. Van Dyk, IPAC, California Institute of Technology; C. K. Lacey,

Naval Research Laboratory; R. A. Sramek, National Radio Astronomy Observatory; and K. W. Weiler, Naval Research Laboratory, report the detection of radio emission from the type-IIn supernova 1998S (IAUC 6829, 6830) with the Very Large Array: “A radio-flux density of 0.954 ± 0.055 mJy at 8.435 GHz (wavelength 3.6 cm) and 1.098 ± 0.075 mJy at 4.860 GHz (wavelength 6.2 cm), both on Oct. 28 in BnA configuration, was measured at the following position: α = 11^{h}46^{m}06^{s}.140, δ = +47^{o}28′55′′.45 (equinox 2000.0; ± 0′′.2 in each coordinate). This is in close agreement with the reported optical position (end figures 06^{s}.18, 55′′.5; IAUC 6831). Previous observations indicate that the supernova was not radio detectable at 8.435 GHz on 1999 June 17, to a limit of < 0.180 mJy (3σ). A list of radio observations, including non-detection upper limits, is available at http://rsd-

www.nrl.navy.mil/7214/weiler/sne-home.html. Filippenko and Moran

(IAUC 6830) predicted radio and x-ray emission from SN 1998S due to spectral evidence, resembling that of the type-IIn SN 1997eg, for strong interaction between the ejecta and circumstellar material. Although the stage of radio evolution is yet to be determined, at an age of ∼ 600 days on Oct. 28, the 6-cm spectral luminosity (for a distance of 17 Mpc; Tully 1988, Nearby Galaxies Catalogue) for SN 1998S of 4 × 10^{26} erg s^{−1} Hz^{−1} is already ∼ 3 times brighter than the well-studied type-IIL SN 1980K at its 6-cm peak flux density. For comparison with other, well-studied type-IIn radio supernovae, SN 1998S is ∼ 5 (SN 1988Z) to 10 (SN 1986J) times less luminous at 6 cm at a similar age. Neither SN 1988Z nor SN 1986J peaked at 6 cm until well after 1000 days of age, so that SN 1998S may still be brightening at cm wavelengths. Detectable radio emission has been found to be closely related to x-ray and late-time optical detectability, so that further multi-wavelength observations are encouraged.”

1999 November 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green